# SLM-Mod: Small Language Models Surpass LLMs at Content Moderation

### v1

@@ Abstract
Large language models (LLMs) have shown promise in many natural language understanding tasks, including content moderation. However, these models can be expensive to query in real-time and do not allow for a community-specific approach to content moderation. To address these challenges, we explore the use of open-source small language models (SLMs) for community-specific content moderation tasks. We fine-tune and evaluate SLMs (less than 15B parameters) by comparing their performance against much larger open- and closed-sourced models. Using 150K comments from 15 popular Reddit communities, we find that SLMs outperform LLMs at content moderation—11.5% higher accuracy and 25.7% higher recall on average across all communities. We further show the promise of cross-community content moderation, which has implications for new communities and the development of cross-platform moderation techniques. Finally, we outline directions for future work on language model based content moderation. [^1] Code & links to HuggingFace models can be found at https://github.com/AGoyal0512/SLM-Mod.
@@ 1 Introduction
This work explores the potential for using SLMs in content moderation, evaluating their performance relative to LLMs in terms of accuracy, recall, and precision across multiple online communities. In particular, we focus on whether fine-tuned SLMs can offer a more resource-efficient approach without compromising on moderation quality. We aim to determine if SLMs can strike a balance between cost-effectiveness and moderation accuracy, ultimately providing a viable alternative to the computationally expensive LLMs for large-scale, community-specific moderation.
Findings.
Our findings reveal that SLMs outperform LLMs at in-domain content moderation tasks having both higher accuracy and recall, with slightly lower precision. We present a case study of r/changemyview and conduct an error analysis on false positives and false negatives to identify trade-offs when using SLMs over LLMs. Next, we highlight the potential of SLMs for cross-domain content moderation tasks and investigate possible reasons for their high performance. Finally we discuss the implications of our findings for content moderation and highlight future directions for improving cross-community approaches to moderation as well as dynamically adapting community-specific models over time.
@@ 2 Related Work
Automated Approaches to Content Moderation
Human moderation is limited in scalability, especially as platforms like Facebook, YouTube, and Reddit experience exponential growth in content volume. To address these limitations, automated approaches have been increasingly adopted. Automated approaches typically use natural language processing and machine learning techniques to flag potentially harmful content for human review Chandrasekharan et al. (2019). For example, n-gram models and sentiment analysis techniques have been used as classic automated approaches to classify content as toxic or non-toxic Davidson et al. (2017); Vigna et al. (2017); Warner and Hirschberg (2012). These methods rely on identifying patterns of word usage and sentiment to make judgments about the harmfulness of content, providing a foundational approach for content moderation. More advanced methods, including deep learning models networks to automatically learn multi-layers of abstract features from raw data have been explored for detecting offensive content Nobata et al. (2016); Badjatiya et al. (2017); Zhang and Luo (2018). Recently, Jha et al. (2024) proposed an LLM and VLM-based framework for online content moderation via meme interventions, and Maity et al. (2023) developed a generative framework for explainable cyber-bullying detection.
LLM-Assisted Content Moderation
@@ 3.2 Models and Configuration
For our study, we evaluate the $4-$bit quantized versions of three small language models (SLMs): Llama-3.1-8b Dubey et al. (2024), Gemma-2-9b Team et al. (2024), and Mistral-nemo-instruct Mistral AI (2024), and three large language models (LLMs): Cohere’s Command R+ Cohere For AI (2024), OpenAI’s GPT-4o; and GPT-4o-mini OpenAI (2024). For the LLMs, a temperature of 0 was used to ensure consistency across runs, and a top_p of $0.75$ was used.
@@ 3.3 Task
The task for the language models is to determine a moderation outcome for a given comment when provided with the comment, its preceding context, and subreddit rules. For a language model $\mathcal{M}$ and a comment $\mathcal{T}$ along with its context $\mathcal{C}$, and subreddit rules $\mathcal{R},$ we prompt the model with a prompt $p.$ Consequently, the moderation task is:
$\mathcal{D}_{\mathcal{M}}\leftarrow\mathcal{M}_{p}(\mathcal{T},\mathcal{C},\mathcal{R})$ (1)
@@
We define two sets of tasks: (a) In-domain Moderation, which involves moderating comments from the same subreddit on which the model was fine-tuned; (b) Cross-domain Moderation, which involves moderating comments from a subreddit different from the one on which the model was fine-tuned. Note that this distinction is primarily useful for SLMs. For off-the-shelf LLMs, the prompt is subreddit-specific but the models themselves are not, so we always prompt with the correct rules.
@@ 3.5 Evaluation Metrics
For our evaluation tasks, we focus on the metrics of accuracy, precision, and recall. These give us a holistic picture of the efficacy of different models on content moderation tasks, along with an insight into how different models handle violating comments in terms of the precision/recall trade-off.
@@ 4.1 Performance Comparison
In this section, we compare the performance of SLMs versus LLMs on in-domain content moderation tasks. Figure 2 provides a visual depiction of the accuracy, recall, and precision across models.
Accuracy
We evaluate the content moderation performance of each model by comparing it to human moderators’ decisions. We find that the fine-tuned SLMs outperform the LLMs by an average 11.5% in accuracy across 15 subreddits. Among the fine-tuned SLMs, Mistral-Nemo-instruct, Gemma-2-9b, and Llama-3.1-8b show accuracies of 77.87%, 77.2%, and 72.5% respectively. Among LLMs, GPT-4o and GPT-4o-mini are the highest-performing models, both with an average accuracy of 65.8%. SLMs’ performance is consistently superior, exhibiting higher accuracy than the LLMs across all subreddits. The highest and lowest average accuracy is achieved by the SLMs on r/changemyview (90.97%) and r/AskReddit (60.4%), and by the LLMs on r/movies (77.4%) and r/askscience (53.4%).
Precision/Recall Trade-off
@@ 4.2 Error Analysis: A case study on r/changemyview
Impact of Content Length
@@
Impact of Content Topic
@@ 5.1 Performance Comparison
In this section, we investigate the cross-domain performance of SLMs on moderation tasks. Figure 4 shows a visual depiction of the findings. We choose three communities examined in prior work Kumar et al. (2024), namely r/IAmA, r/askscience, and r/movies as our three subreddits for the discussion of cross-domain performance. Appendix D provides complete cross-domain results.
@@
Furthermore, we see that there are some models that show a promising overall cross-domain performance. Specifically, models trained on r/changemyview, r/nba, and r/movies show an average cross-domain accuracy between 64.6 and 67.8%, and can therefore be used as ‘meta-experts’ to moderate communities that would benefit more from a cross-domain moderation as compared to in-domain moderation. One example is r/AskReddit, which would benefit from the cross-domain models from 10 other communities compared to its in-domain model based on Llama-3.1-8b and Gemma-2-9b, and 7 others for Mistral-NeMo-Instruct.
@@ 6.1 Shift from generalist LLMs to specialist SLMs for content moderation
Our findings indicate a significant advantage in employing SLMs for content moderation tasks over LLMs, consistently showing superior performance in terms of both accuracy and recall. Our findings suggest that SLMs adopt a more aggressive approach compared to LLMs, resulting in significantly higher recall for SLMs, though at the cost of slightly lower precision relative to LLMs. However, prior work has shown that moderators, who are overburdened already, may not be able to attend to all instances of norm-violating and undesirable behavior Park et al. (2022); Chandrasekharan et al. (2019). Hence, the higher recall of SLMs compared to LLMs could actually provide a benefit for more reliable detection of harmful content that could otherwise be overlooked and stay on the platform for longer—potentially leading to further undesirable outcomes Lambert et al. (2022).
@@ 6.4 Instability of Closed-source LLMs
Given the black-box nature of closed-source models and the lack of knowledge of upcoming updates, ensuring consistent moderation performance would pose a challenge when using closed-source models. Moderators may instead consider using open-source language models, which tend to be more stable since moderators can choose not to update the models and can be fine-tuned more easily to meet specific moderation requirements.
@@ 7 Conclusion
This paper examined the effectiveness of small language models (SLMs) and large language models (LLMs) in content moderation tasks. We found that SLMs with less than 15b parameters, such as Gemma-2-9b and Mistral-Nemo-Instruct, consistently outperformed significantly larger LLMs like GPT-4o at identifying undesirable content. SLMs showed a higher recall, indicating their superior ability to flag a wider range of undesirable content, which can be crucial for effective moderation on large-scale platforms. In addition to in-domain moderation tasks, we uncovered the potential of SLMs to be used in cross-community moderation tasks. Cross-community moderation can benefit smaller communities with fewer resources to train their own in-domain models or newer communities where community norms are still emerging. We provided qualitative insights into the tradeoffs arising from content length and topic that could impact SLMs and LLMs differently. Overall, SLMs offer an effective, scalable, and cost-effective solution for content moderation, achieving a strong balance between accuracy, recall, and stability. Future work can further improve cross-domain performance by training models that learn shared notions of norms and values, and explore hybrid models that leverage complementary strengths of SLMs and LLMs to enhance moderation capabilities across diverse online communities and platforms.
@@ Acknowledgments
A.G. was supported by compute credits from a Cohere For AI Research Grant. This work used the Delta system at the National Center for Supercomputing Applications through allocation #240481 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ Appendix A Subreddits Statistics
Table 2: Subreddit subscriber sizes and descriptions
@@
Table 3: Model Accuracy by Subreddit
Subreddit llama-3.1-8b gemma-2-9b mistral-nemo-instruct-12b Command-R-Plus gpt-4o-mini gpt-4o gpt-4o-2024-05-13 gpt3.5-turbo
@@
Table 4: Model Precision by Subreddit
Subreddit llama-3.1-8b gemma-2-9b mistral-nemo-instruct-12b Command-R-Plus gpt-4o-mini gpt-4o gpt-4o-2024-05-13 gpt3.5-turbo
@@
Table 5: Model Recall by Subreddit
Subreddit llama-3.1-8b gemma-2-9b mistral-nemo-instruct-12b Command-R-Plus gpt-4o-mini gpt-4o gpt-4o-2024-05-13 gpt3.5-turbo

### v2

@@ Abstract
Large language models (LLMs) have shown promise in many natural language understanding tasks, including content moderation. However, these models can be expensive to query in real-time and do not allow for a community-specific approach to content moderation. To address these challenges, we explore the use of open-source small language models (SLMs) for community-specific content moderation tasks. We fine-tune and evaluate SLMs (less than 15B parameters) by comparing their performance against much larger open- and closed-sourced models in both a zero-shot and few-shot setting. Using 150K comments from 15 popular Reddit communities, we find that SLMs outperform zero-shot LLMs at content moderation—11.5% higher accuracy and 25.7% higher recall on average across all communities. Moreover, few-shot in-context learning leads to only a marginal increase in the performance of LLMs, still lacking compared to SLMs. We further show the promise of cross-community content moderation, which has implications for new communities and the development of cross-platform moderation techniques. Finally, we outline directions for future work on language model based content moderation. [^1] Code: https://github.com/AGoyal0512/SLM-Mod.
@@ 1 Introduction
This work explores the potential for using SLMs in content moderation, evaluating their performance relative to LLMs in terms of accuracy, recall, and precision across multiple online communities on Reddit. In particular, we focus on whether fine-tuned SLMs can offer a more resource-efficient approach without compromising on moderation quality. We aim to determine if SLMs can strike a balance between cost-effectiveness and moderation accuracy, ultimately providing a viable alternative to the computationally expensive LLMs for large-scale, community-specific moderation.
Findings:
Our findings reveal that fine-tuned SLMs outperform zero-shot LLMs at in-domain content moderation tasks having both higher accuracy and recall, with slightly lower precision. We present a case study of r/changemyview and conduct an error analysis on false positives and false negatives to identify trade-offs when using SLMs over LLMs. We show that even under a few-shot setting, LLMs lack performance compared to SLMs. Moreover, SLMs have a higher AUC score compared to LLMs on more realistic imbalanced datasets. Next, we highlight the potential of SLMs for cross-domain content moderation tasks and investigate possible reasons for their high performance. Finally we discuss the implications of our findings for content moderation and highlight future directions for improving cross-community approaches to moderation as well as dynamically adapting community-specific models over time.
@@ 2 Related Work
Automated Approaches to Content Moderation:
Due to the limited scalability of human moderation, automated approaches have been increasingly adopted. Automated approaches typically use natural language processing and machine learning techniques to flag potentially harmful content for human review Chandrasekharan et al. (2019). For example, n-gram models and sentiment analysis techniques have been used as classic automated approaches to classify content as toxic or non-toxic Davidson et al. (2017); Vigna et al. (2017); Warner and Hirschberg (2012). These methods rely on identifying patterns of word usage and sentiment to make judgments about the harmfulness of content, providing a foundational approach for content moderation. More advanced methods, including deep learning models to automatically learn multi-layers of abstract features from raw data have been explored for detecting offensive content Nobata et al. (2016); Badjatiya et al. (2017); Zhang and Luo (2018). Recently, Jha et al. (2024) proposed an LLM and VLM-based framework for online content moderation via meme interventions, and Maity et al. (2023) developed a generative framework for explainable cyber-bullying detection.
LLM-Assisted Content Moderation:
@@ 3.2 Models and Configuration
For our study, we evaluate the $4-$bit quantized versions of three small language models (SLMs): Llama-3.1-8b Dubey et al. (2024), Gemma-2-9b Team et al. (2024), and Mistral-nemo-instruct Mistral AI (2024), and three large language models (LLMs): Cohere’s Command R+ Cohere For AI (2024), OpenAI’s GPT-4o; and GPT-4o-mini OpenAI (2024). For the LLMs, a temperature of 00 was used to ensure consistency across runs, and a top_p of $0.75$ was used.
@@ 3.3 Task
The task for the language models is to determine a moderation outcome for a given comment when provided with the comment, its preceding context, and subreddit rules. For a language model $\mathcal{M}$ and a comment $\mathcal{T}$ along with its context $\mathcal{C}$, and subreddit rules $\mathcal{R},$ we prompt the model with a prompt $p_{k}$ where $k\in\{0,2,4\}$ represents the number of in-context examples provided to the model. Consequently, the moderation task is:
$\mathcal{D}_{\mathcal{M}}\leftarrow\mathcal{M}_{p_{k}}(\mathcal{T},\mathcal{C},\mathcal{R})$ (1)
@@
We define two sets of tasks: (a) In-domain Moderation, which involves moderating comments from the same subreddit on which the model was fine-tuned; (b) Cross-domain Moderation, which involves moderating comments from a subreddit different from the one on which the model was fine-tuned. Note that this distinction is primarily useful for SLMs. For off-the-shelf LLMs, the prompt is subreddit-specific but the models themselves are not, so we always prompt with the correct rules. Furthermore, for the LLMs we use both a zero-shot and few-shot evaluation setup to provide off-the-shelf LLM with some contextual knowledge about the content moderation task.
3.4 Evaluation Data Distributions
In real-world moderation settings, the distribution of moderated and unmoderated comments is often heavily skewed. Platforms typically do not report the rates of comments moderated as it ties in with the public perception of the platform. However, Park et al. (2022a) estimated the prevalence of anti-social behavior in 97 popular subreddits at around 5% (6.25% in 2016 and 4.28% in 2020). As a result, we evaluate the performance of models on two kind of data distributions: (a) Balanced, where the number of moderated and unmoderated cases in the test-set are equal; (b) Imbalanced, where the number of moderated comments is 1%, 5%, or 10% of the total number of comments in the test-set.
@@ 3.6 Evaluation Metrics
For our evaluation tasks in the balanced setting, we focus on the metrics of accuracy, precision, and recall. These give us a holistic picture of the efficacy of different models on content moderation tasks, along with an insight into how different models handle violating comments in terms of the precision/recall trade-off. However for the imbalanced setting, we focus on the AUC score as it is a better representation of model performance being less sensitive to class distribution imbalances.
@@ 4.1 Performance Comparison
In this section, we compare the performance of fine-tuned SLMs versus zero-shot LLMs on in-domain content moderation tasks in a balanced evaluation setting. Figure 2 provides a visual depiction of the accuracy, recall, and precision across models.
Accuracy:
We evaluate the content moderation performance of each model by comparing it to human moderators’ decisions. We find that the fine-tuned SLMs [^3] For detailed performance of base SLMs, see Appendix E. outperform the LLMs by an average 11.5% in accuracy across 15 subreddits. Among the fine-tuned SLMs, Mistral-NeMo-instruct, Gemma-2-9b, and Llama-3.1-8b show accuracies of 77.87%, 77.2%, and 72.5% respectively. Among LLMs, GPT-4o and GPT-4o-mini are the highest-performing models, both with an average accuracy of 65.8%. SLMs’ performance is consistently superior, exhibiting higher accuracy than the LLMs across all subreddits. The highest and lowest average accuracy is achieved by the SLMs on r/changemyview (90.97%) and r/AskReddit (60.4%), and by the LLMs on r/movies (77.4%) and r/askscience (53.4%).
Precision/Recall Trade-off:
@@ 4.2 Error Analysis: A case study on r/changemyview
Impact of Content Length:
@@
Impact of Content Topic:
@@
4.3 Does In-Context Learning Improve the Performance of LLMs?
Table 1: In-Context Learning Performance of LLMs: LLM performance comparison across subreddits for 2– and 4–shot ICL. Colored by improvement over 0–shot LLM baselines in Section 4.1: Reduction, Improvement by $<4\%$, Improvement by $\geq 4\%$.
Subreddit \ Model GPT-4o-mini GPT-4o Command R+
n–shot $n=2$ $n=4$ $n=2$ $n=4$ $n=2$ $n=4$
r/nba 73.1 74.2 75.2 75.8 73.3 74.7
r/aww 65.7 70.8 70.7 70.2 70.1 70.3
r/movies 71.7 77.2 80.8 79.1 77.8 76.9
r/politics 62.1 69.8 74.7 75.8 69.8 70.3
r/changemyview 68.8 71.8 75.5 76.2 74.6 71.1
Since we have used LLMs in a zero-shot manner so far, we investigated whether providing examples to the LLM for in-context learning (ICL) could improve their performance. To do so, we use $n-$shot ICL ($n\in\{2,4\}$) and compare to the zero-shot setup in Table 1 for the five subreddits where the LLMs showed the highest performance. In order to mitigate potential biases in the chosen examples for ICL, we randomly sampled $2$ or $4$ examples from the training split for each setting.
Figure 4: Imbalanced Distribution Evaluation. Best performing SLM (Mistral-NeMo-Instruct) and LLM (GPT-4o) on $1\%$, $5\%$, and $10\%$ imbalance-level test split of r/changemyview and r/AskReddit by AUC scores. Error bars represent standard deviation over 30 seeds.
We see that ICL either leads to a reduction in performance of the LLMs or provides marginal gains ($<4\%$) for both GPT-4o-mini and GPT-4o. For Command R+ we see a performance gain of over $4\%$ in most cases, but is still outperformed by GPT-4o. Overall, we observe that including ICL examples for LLMs in content moderation task can be unstable and fails to help LLMs match the performance of SLMs. Our results match those of Guo et al. (2023) on the related task of hate-speech detection where few-shot prompting leads to the lowest performance across various prompting techniques used by the researchers.
4.4 How Does Imbalanced Data Affect the Performance of SLMs versus LLMs?
We now evaluate the content moderation performance of language models on imbalanced datasets, as balanced sampling results in a less challenging distribution than real-time moderation data where the number of removals are typically low ($<10\%$).
We compute the test metrics at different imbalance thresholds: 1%, 5%, and 10% moderated comments and remaining unmoderated comments. Due to the possibility of high variance based on the sample, we collate AUC scores for 30 different runs. We report results here for 2 subreddits—r/changemyview and r/AskReddit—where the SLMs had the best and the worst performance, respectively, in the balanced setting. We pick the SLM and LLM with the best performance at each imbalance level, which were Mistral-NeMo-Instruct and GPT-4o in all cases.
From Figure 4 we notice that even under an imbalanced distribution, Mistral-Nemo-Instruct outperforms GPT-4o on moderating content from both subreddits with average AUC of 0.7915, 0.785, and 0.781 compared to the average AUC of 0.698, 0.687, 0.683 of GPT-4o at a 1%, 5%, and 10% imbalance level respecively.
@@ 5.1 Performance Comparison
In this section, we investigate the cross-domain performance of SLMs on moderation tasks. Figure 5 shows a visual depiction of the findings. We choose three communities examined in prior work, namely r/IAmA, r/askscience, and r/movies as our three subreddits for the discussion of cross-domain performance Kumar et al. (2024)Appendix D provides complete cross-domain results.
@@
Furthermore, we see that there are some models that show a promising overall cross-domain performance. Specifically, models trained on r/changemyview, r/nba, and r/movies show an average cross-domain accuracy between 64.6% and 67.8%, and can therefore be used as ‘meta-experts’ to moderate communities that would benefit more from a cross-domain moderation as compared to in-domain moderation. One example is r/AskReddit, which would benefit from the cross-domain models from 10 other communities compared to its in-domain model based on Llama-3.1-8b and Gemma-2-9b, and 7 others for Mistral-NeMo-Instruct.
@@ 6.1 Shift from generalist LLMs to specialist SLMs for content moderation
Our findings indicate a significant advantage in employing SLMs for content moderation tasks over LLMs, consistently showing superior performance in terms of both accuracy and recall. Our findings suggest that SLMs adopt a more aggressive approach compared to LLMs, resulting in significantly higher recall for SLMs, though at the cost of slightly lower precision relative to LLMs. However, prior work has shown that moderators, who are overburdened already, may not be able to attend to all instances of norm-violating and undesirable behavior Park et al. (2022b); Chandrasekharan et al. (2019). Hence, the higher recall of SLMs compared to LLMs could actually provide a benefit for more reliable detection of harmful content that could otherwise be overlooked and stay on the platform for longer—potentially leading to further undesirable outcomes Lambert et al. (2022).
@@ 6.4 Instability of Closed-source LLMs
Given this black-box nature of closed-source models, ensuring consistent moderation performance would pose a challenge when using closed-source models. Moderators may instead consider using open-source language models, which tend to be more stable since moderators can choose not to update the models and can be fine-tuned more easily to meet specific moderation requirements.
@@ 7 Conclusion
This paper examined the effectiveness of small language models (SLMs) and large language models (LLMs) in content moderation tasks. We found that SLMs with less than 15b parameters, such as Gemma-2-9b and Mistral-Nemo-Instruct, consistently outperformed significantly larger LLMs like GPT-4o at identifying undesirable content. SLMs showed a higher recall, indicating their superior ability to flag a wider range of undesirable content, which can be crucial for effective moderation on large-scale platforms. We also found that SLMs have higher AUC scores than LLMs under realistic imbalanced data conditions, and that even with in-context examples LLMs fail to match SLMs. In addition to in-domain moderation tasks, we uncovered the potential of SLMs to be used in cross-community moderation tasks. Cross-community moderation can benefit smaller communities with fewer resources to train their own in-domain models or newer communities where community norms are still emerging. We provided qualitative insights into the trade-offs arising from content length and topic that could impact SLMs and LLMs differently. Overall, SLMs offer an effective, scalable, and cost-effective solution for content moderation, achieving a strong balance between accuracy, recall, and stability. Future work can further improve cross-domain performance by training models that learn shared notions of norms and values, and explore frameworks that leverage complementary strengths of SLMs and LLMs to enhance moderation capabilities across diverse online communities.
@@ Acknowledgments
The authors sincerely thank the anonymous reviewers for their constructive feedback on our work during the ARR peer review stage. A.G. was supported by compute credits from a Cohere For AI Research Grant. This work used the Delta system at the National Center for Supercomputing Applications through allocation #240481 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ Appendix A Subreddits Statistics
Table 3: Base Small Language Model (SLM) Metrics by Subreddit.
llama-3.1-8b gemma-2-9b mistral-nemo-instruct
Subreddit Precision Recall Accuracy Precision Recall Accuracy Precision Recall Accuracy
r/AskHistorians 0.511 0.594 0.513 0.000 0.000 0.500 0.772 0.044 0.516
r/AskReddit 0.514 0.622 0.516 0.846 0.011 0.504 0.701 0.286 0.582
r/AskWomen 0.498 0.600 0.497 0.667 0.004 0.501 0.820 0.201 0.578
r/Futurology 0.507 0.605 0.508 0.500 0.005 0.500 0.800 0.308 0.616
r/Games 0.495 0.634 0.494 0.656 0.040 0.509 0.868 0.132 0.556
r/IAmA 0.398 0.654 0.460 0.353 0.009 0.592 0.816 0.410 0.724
r/anime 0.500 0.652 0.500 0.500 0.004 0.500 0.819 0.068 0.526
r/askscience 0.497 0.604 0.496 0.167 0.001 0.498 0.744 0.067 0.522
r/aww 0.510 0.622 0.512 0.333 0.002 0.499 0.875 0.371 0.659
r/changemyview 0.512 0.661 0.515 0.400 0.004 0.499 0.861 0.408 0.671
r/history 0.497 0.527 0.496 0.567 0.017 0.502 0.802 0.097 0.536
r/movies 0.490 0.630 0.486 0.556 0.005 0.500 0.912 0.560 0.753
r/nba 0.491 0.623 0.489 0.320 0.008 0.496 0.892 0.531 0.734
r/politics 0.512 0.577 0.514 1.000 0.001 0.500 0.817 0.401 0.656
r/science 0.510 0.535 0.510 0.295 0.013 0.491 0.723 0.191 0.559
Table 4: Subscriber sizes and descriptions of the 15 subreddits studied in this work.
@@
Table 5: Fine-tuned Small Language Model (SLM) and Large Language Model (LLM) Accuracy by Subreddit.
Subreddit llama-3.1-8b gemma-2-9b mistral-nemo-instruct Command-R-Plus gpt-4o-mini gpt-4o gpt-4o-2024-05-13 gpt3.5-turbo
@@
Table 6: Fine-tuned Small Language Model (SLM) and Large Language Model (LLM) Precision by Subreddit.
Subreddit llama-3.1-8b gemma-2-9b mistral-nemo-instruct Command-R-Plus gpt-4o-mini gpt-4o gpt-4o-2024-05-13 gpt3.5-turbo
@@
Table 7: Fine-tuned Small Language Model (SLM) and Large Language Model (LLM) Recall by Subreddit.
Subreddit llama-3.1-8b gemma-2-9b mistral-nemo-instruct Command-R-Plus gpt-4o-mini gpt-4o gpt-4o-2024-05-13 gpt3.5-turbo